# A determinantal formula for cluster variables in cluster algebras from surfaces

### v1

@@ 1 Introduction
Cluster algebras are a broad class of commutative algebras introduced by Fomin and Zelevinsky in [5] as a framework for the study of total positivity. They are recursively defined from an initial set of data by a combinatorial process called mutation. In [4], Fomin, Shapiro and Thurston introduced and studied a class of cluster algebras whose combinatorics are governed by flips of (tagged) triangulations in surfaces with marked points. In [8], Musiker, Schiffler and Williams gave a cluster expansion formula for cluster variables in these cluster algebras in terms of the perfect matchings of certain planar graphs called snake graphs, see Theorem 3.1. These graphs were studied further in [2], where a surprising connection to continued fractions was developed. In this paper, we apply a classical idea in matching theory to provide a reformulation of Musiker’s, Schiffler’s and Williams’ cluster expansion formula into a determinantal formula. This provides a way to circumvent the enumeration of the perfect matchings of the relevant snake graph.
@@
We begin by briefly recalling the definition of cluster algebras (with trivial coefficients) obtained from surfaces, following [4]. Let $\Sigma$ be a Riemann surface with boundary $\partial\Sigma$, and let $\mathbb{M}\subset\Sigma$ be a non empty finite set of marked points such that $\mathbb{M}\cap C\neq\emptyset$ for each connected component $C$ of $\partial\Sigma$. We will call the pair $(\Sigma,\mathbb{M})$ a bordered surface. Call the marked points in $\partial\Sigma$ ideal points, and those in the interior of $\Sigma$ are punctures. A boundary arc of $\Sigma$ is a curve contained in $\delta\Sigma$ with endpoints in $M$. An arc in $\Sigma$ is a curve $\gamma$ with ideal points as endpoints, no self intersections, and not homotopic to a boundary arc or to a point; similarly, a All such curves are considered up to homotopy. Two arcs are compatible if they do not intersect in their relative interiors, up to homotopy. A maximal collection $T$ of pairwise compatible arcs in $\Sigma$ along with all of the boundary arcs of $\Sigma$ is an ideal triangulation of $(\Sigma,\mathbb{M})$. All triangulations of $(\Sigma,\mathbb{M})$ have $66g+3b+3p+c-$ 6 arcs.
@@
Fix a surface with marked points $(\Sigma,\mathbb{M})$ and an ideal triangulation $T_{0}$ of $(\Sigma,\mathbb{M})$ with valency at least 3 at each puncture with arcs $\tau_{1},\ldots,\tau_{n}$. For each triangle $\Delta$ in $T_{0}$, let $B^{\Delta}=(b_{i,j}^{\Delta})_{1\leq i\leq n,1\leq j\leq n}$, where
$b_{i,j}^{\Delta}=\begin{cases}1&\text{if $t_{j}$ follows $t_{i}$ in clockwise order,}\\ -1&\text{if $t_{j}$ follows $t_{i}$ in counterclockwise order,}\\ 0&\text{otherwise}.\end{cases}$
@@
The signed adjacency matrix $B=B(T_{0})$ is given by $B(T_{0})=\sum_{\Delta}B^{\Delta}$, where the sum is over all non self-folded triangles. Take a set of $n$ algebraically independent rational functions, one for each arc in $T_{0}$, $x_{0}=\{x_{\tau_{1}},\ldots,x_{\tau_{n}}\}$, that generate the field $\mathbb{C}(u_{1},\ldots,u_{n})$. Such a set is called an initial cluster. The pair $(B(T_{0}),x_{0})$ is called an initial seed. For each $k\in\{1,\ldots,n\}$ define the mutation in direction k by the following rule. Let $\mu_{k}(x_{\tau_{i}})=x_{\tau_{i}}$ if $i\neq k$, and
$\mu_{k}(x_{\tau_{k}})=\frac{1}{x_{\tau_{k}}}\left(\prod_{B_{k,j}>0}x_{\tau_{j}}^{B_{k,j}}+\prod_{B_{k,j}<0}x_{\tau_{j}}^{-B_{k,j}}\right)\,,$
so that $\mu_{k}(x_{0})=x_{0}\setminus\{x_{k}\}\cup\{\mu_{k}(x_{k})\}$. Define
$\mu_{k}(B)_{i,j}=\begin{cases}-B_{i,j}&\text{if }k\in\{i,j\}\\ B_{i,j}+\frac{|B_{i,k}|B_{k,j}+B_{i,k}|B_{k,j}|}{2}&\text{otherwise}\end{cases}$
@@
Write $\mu_{k}(B(T_{0}),x_{0}):=(\mu_{k}(B(T_{0})),\mu_{k}(x_{0}))$. Observe that $\mu_{k}(B(T_{0}))$ is the adjacency matrix of the triangulation $T^{\prime}$ obtained by flipping the edge $\tau_{k}$. The set $\mu_{k}(x_{0})$ is again a cluster. Let $\mathcal{X}$ be the set of cluster variables; that is, all variables contained in all possible clusters obtained by sequences of mutations starting from $x_{0}$. The cluster algebra (with trivial coefficients) associated to the surface with marked points $(\Sigma,\mathbb{M})$ is the algebra $\mathcal{A}=\mathcal{A}(T_{0},x_{0})=\mathbb{C}[\mathcal{X}]$. It is independent of the chosen triangulation $T_{0}$, so it really is associated to the surface $(\Sigma,\mathbb{M})$. Furthermore, there is a biyective correspondence between clusters of $\mathcal{A}$ and triangulations of $(\Sigma,\mathbb{M})$, as well as between cluster variables and arcs in $\Sigma$.
@@ 2 Perfect matchings
All graphs considered in this paper will be finite and loopless. Let $\mathcal{G}$ be such a graph with vertex set $V=\{v_{1},\ldots,v_{n}\}$ and edge set $E$. Denote by $E(i,j)$ the set of edges between $v_{i}$ and $v_{j}$. Write $s(G)$ for the underlying simple graph of $G$, and $m(i,j)=|E(i,j)|$. Of course, if $G$ is simple then $s(G)=G$, and $m(i,j)\in\{0,1\}$ for all $1\leq i,j\leq n$. If $\overrightarrow{G}$ is an orientation of $G$, then $E(i,j)$ denotes the set of arrows with tail $v_{i}$ and head $v_{j}$.
A perfect matching of $G$ is a subset $M\subseteq E$ such that each $v\in V$ is an endpoint of exactly one member of $M$. Thus, $|M|=|V|/2$, so a graph that admits a perfect matching must have an even number of vertices. Call such a graph even. Denote by $\operatorname{Match}(G)$ the set of perfect matchings of $G$, and let $\Phi(G):=|\operatorname{Match}(G)|.$
We will use a simple but useful device to keep track of all perfect matchings of $G$, the so-called method of variables as in [7].
@@
(Weighted adjacency matrix) Let $G$ be a graph with vertex set $V=\{v_{1},\ldots,v_{n}\}$. For each edge $e$ of $s(G)$, consider the formal variable $z_{e}$, and let $a_{i,j}=m(i,j)z_{e}$. The matrix $A_{G}(z):=(a_{i,j})$ is called the weighted adjacency matrix of $G$. Furthermore, if $\overrightarrow{G}$ is an orientation of $G$ with no oriented 2-cycles, write
$b_{i,j}:=\epsilon(i,j)m(i,j)z_{e}\,,\quad\text{where}\quad\epsilon(i,j):=\begin{cases}1&E(i,j)\neq\emptyset\\ -1&E(j,i)\neq\emptyset\\ 0&\text{ otherwise}\end{cases}\,.$
$B_{\overrightarrow{G}}(z):=(b_{i,j})$ is the weighted skew-symmetric adjacency matrix of $\overrightarrow{G}$.
@@
The matrices of definition 2.1 are defined up to a relabeling of the vertex set of $G$, which corresponds to a simultaneous permutation of their rows and columns. By setting each variable $z_{e}=1$, we recover the usual adjacency matrices of the graphs $G$ and $\overrightarrow{G}$, so we denote them by $A_{G}(1)$ and $B_{\overrightarrow{G}}(1)$, respectively.
@@
$\phi_{G}:=\sum_{\{e_{1},\ldots,e_{k}\}\in\operatorname{Match}(G)}z_{e_{1}}\cdots z_{e_{k}}$
@@
The term $c_{P}$ is independent of the order of labeling of each part in $P$. Indeed, switching the roles of $i_{k}$ and $j_{k}$ changes the sign of $\sigma_{P}$ by $-1$, but $c_{j_{k},i_{k}}=-c_{i_{k},j_{k}}$. Thus, $c_{P}$ really does depend only on the partition $P$. Note that such a partition corresponds to a perfect matching of $K_{2n}$, the complete graph in $2n$ vertices, and that $c_{P}\neq 0$ if and only if such a potential matching corresponds to an actual matching of $G$.
@@
Let $G$ be an graph with an even number of vertices, and let $\overrightarrow{G}$ be an arbitrary orientation of $G$. The sum of the absolute value of the coefficients of $\operatorname{pff}(B_{\overrightarrow{G}}(z))$ is $\Phi(G)$.
@@
Take the vertex set $V$ of $G$ to be $\{1,\ldots,n\}$, and consider the simple graph $s(G)$ with edge set $E^{\prime}$. Let $k=n/2$. Each partition $P=\{\{i_{l},j_{l}\}\mid 1\leq l\leq k\}$ as in definition 2.3 corresponds to a possible set of independent edges in the simple graph $s(G)$. For $C=B_{\overrightarrow{G}}(z)$, the associated term $c_{P}$ is
@@
As Tutte first observed [9], the previous argument and Lemma 2.1 imply that a graph $G$ admits a perfect matching if and only if there is some orientation $\overrightarrow{G}$ for which $\det(B_{\overrightarrow{G}}(z))\neq 0$. Call the polynomial $\operatorname{pff}(B_{\overrightarrow{G}}(z))$ the pfaffian of $\overrightarrow{G}$, and denote it by $\operatorname{pff}(\overrightarrow{G})$.
@@
Let $\overrightarrow{G}$ be an oriented graph. Then $|\operatorname{pff}(B_{\overrightarrow{G}}(1))|\leq\Phi(G)$.
An orientation $\overrightarrow{G}$ of $G$ for which the equality is attained in the previous corollary is called a pfaffian orientation. Likewise, a graph is pfaffian if it admits a pfaffian orientation. Observe that for such an orientation, the pfaffian of $\overrightarrow{G}$ is a polynomial whose coefficients all have the same sign. Indeed, in such a case $\operatorname{pff}(\overrightarrow{G})=\pm\phi_{G}$. We cite a useful characterization of pfaffian orientations taken from [7]. Call a cycle $C$ of a graph $G$ nice if $G-V(C)$ admits a perfect matching.
@@
([7], Theorem 8.3.2) Let $G$ be an even graph and $\overrightarrow{G}$ an orientation of $G$. Then $\overrightarrow{G}$ is pfaffian if and only if, when traversing each even nice cycle of $G$ in an arbitrary direction, it has an odd number of arrows that agree with the orientation $\overrightarrow{G}$.
Furthermore, the following theorem is classic, and a proof can be found in [7] where it appears as Theorem 8.3.4.
@@
Let $\overrightarrow{G}$ be bipartite with bipartition $V=U\cup W$ such that $|U|=|W|$ (note that if $\overrightarrow{G}$ has a perfect matching, this is the case). The weighted skew symmetric adjacency matrix of $\overrightarrow{G}$ takes the form
$B_{\overrightarrow{G}}(z)=\begin{pmatrix}0&M_{\overrightarrow{G}}(z)\\ -M_{\overrightarrow{G}}(z)^{T}&0\end{pmatrix}\,,$
for any labeling $V=\{u_{1},\ldots,u_{k},w_{1},\ldots,w_{l}\}$ such that $u_{i}\in U$ and $w_{j}\in W$ for $1\leq i\leq k$, $1\leq j\leq l$. $M_{\overrightarrow{G}}(x)$ is called the weighted biadjacency matrix of $\overrightarrow{G}$. The following lemma is given as an exercise in [6] (though there seems to be a sign missing therein). We include a proof.
@@
Let $M$ be a square matrix of size $n$. If
$B=\begin{pmatrix}0&M\\ -M^{T}&0\end{pmatrix}\,,$
then $\operatorname{pff}(B)=(-1)^{n(n-1)/2}\det M$.
@@
Let $V=\{1,\ldots,2n\}$, $U=\{1,\ldots,n\}$ and $W=\{n+1,\ldots,2n\}$. Fix a partition $P=\{\{i_{1},j_{1}\},\ldots,\{i_{n},j_{n}\}\}$. By remark 2.2, we can suppose that $j_{i}=n+i$ for each $1\leq i\leq n$, so
@@
Note that $i_{1},\ldots,i_{n}\leq n<n+1=j_{1}$, so the indices $i_{l}$ amount to $n-1$ inversions with respect to $j_{1}$. Similarly, $n-2$ of the $i_{l}$ give $n-2$ inversions with $j_{2}$, and so on. This gives a total of $n(n-1)/2$ inversions. Thus,
@@
Thus, $b_{P}=(-1)^{n(n-1)/2}\operatorname{sgn}(\sigma_{P}^{\prime})b_{i_{1},n+1}\cdots b_{i_{n},2n}$. Summing over all partitions $P$, we obtain the result. ∎
In particular, calculating the pfaffian of a bipartite graph amounts to computing the determinant of its biadjacency matrix.

### v2

@@ 1 Introduction
Cluster algebras are a broad class of commutative algebras introduced by Fomin and Zelevinsky in [5] as a combinatorial framework for the study of total positivity and canonical bases. They are recursively defined from an initial set of data by a combinatorial process called mutation. In [4], Fomin, Shapiro, and Thurston introduced and studied a class of cluster algebras whose combinatorics are governed by flips of tagged triangulations in surfaces with marked points. In [8], Musiker, Schiffler and Williams gave cluster expansion formulae for the cluster variables of these cluster algebras in terms of the perfect matchings of certain planar graphs called snake graphs. These graphs were studied further in [2], where a remarkable connection to continued fractions was developed.
In this note, we use a classical idea of matching theory to give a reformulation of Musiker’s, Schiffler’s, and Williams’ cluster expansion formula into a determinantal formula for the cluster variables associated to plain arcs (see Proposition 3.3). This gives a simple way to avoid the enumeration of the perfect matchings of the corresponding snake graph. The formula works by using an algebraic device to count all of the perfect matchings of a graph. The formulae for cluster variables associated to tagged arcs in [8] are given in terms of more restrictive matchings, simplified in [10], were the expansion formulae involve good matchings of so-called loop graphs. Thus the scope of our strategy is limited to the case of plain arcs.
@@
We begin by briefly recalling the definition of surface cluster algebras following [4]. Let $\Sigma$ be a Riemann surface with boundary $\partial\Sigma$, and let $\mathbb{M}\subset\Sigma$ be a non-empty finite set of marked points such that $\mathbb{M}\cap C\neq\emptyset$ for each connected component $C$ of $\partial\Sigma$. The pair $(\Sigma,\mathbb{M})$ is a bordered surface, the marked points are called $\partial\Sigma$ ideal points, and those in the interior of $\Sigma$ punctures. A boundary arc of $\Sigma$ is a curve with endpoints in $M$ that is contained in $\delta\Sigma$, and an arc in $\Sigma$ is a curve $\gamma$ with endpoints in $\mathbb{M}$ such that its relative interior is disjoint from $M$, it has no self-intersections, and it is not homotopic to a boundary arc or to a point. All such arcs are considered up to isotopy. Two arcs are said to be compatible if they do not intersect in their relative interiors, up to homotopy. A maximal collection $T$ of distinct pairwise compatible arcs in $\Sigma$ along with all the boundary arcs of $\Sigma$ is an ideal triangulation of $(\Sigma,\mathbb{M})$. All ideal triangulations of $(\Sigma,\mathbb{M})$ have 6g+3b+3p+c66g+3b+3p+c-6 arcs. It may happen that not all triangles of an ideal triangulation $T$ have three sides: when a loop $l$ encircles a puncture $p$, there is exactly one arc $r$ joining the base of $l$ to $p$ that is compatible with $l$. The shape formed by $l$ and $r$ is called a self-folded triangle, where $r$ is the self-folded side.
@@
Pick an initial ideal triangulation $T_{0}$ with arcs $\tau_{1},\ldots,\tau_{n}$. For each triangle $\Delta$ in $T_{0}$, let $B^{\Delta}=(b_{i,j}^{\Delta})_{1\leq i\leq n,1\leq j\leq n}$, where
[…]
The signed adjacency matrix $B=B(T_{0})$ is given by $B=\sum_{\Delta}B^{\Delta}$, where the sum is taken over all triangles of $T_{0}$ that are not self-folded.
Consider the tropical semifield $\mathbb{P}:=\operatorname{Trop}(y_{1},\ldots,y_{n})$ defined as the multiplicative abelian group freely generated by $y_{1},\ldots,y_{n}$ with the auxiliary sum
$\prod_{i=1}^{n}y_{i}^{a_{i}}\oplus\prod_{i=1}^{n}y_{i}^{b_{i}}:=\prod_{i=1}^{n}y_{i}^{\min(a_{i},b_{i})}\,.$
@@
Take a tuple of $\textbf{x}_{0}=(x_{1},\ldots,x_{n})$ of algebraically independent generators of the field of rational functions $\mathbb{Q}\mathbb{P}(u_{1},\ldots,u_{n})$. Such a tuple is called a cluster. Let $\mathbf{y}_{0}\in\mathbb{P}^{n}$. The triple $(\textbf{x}_{0},\textbf{y}_{0},B)$ is called the initial seed. For each $k\in\{1,\ldots,n\}$ define the mutation in direction k $\mu_{k}(\textbf{x}_{0},\textbf{y}_{0},B_{0}):=(\mathbf{x}^{\prime},\mathbf{y}^{\prime},B^{\prime})$ by the following rule:
  • The matrix $B^{\prime}=(b^{\prime}_{ij})$ is defined by
    $b^{\prime}_{i,j}=\begin{cases}-b_{i,j}&\text{if }k\in\{i,j\}\,,\\ b_{i,j}+\frac{|b_{i,k}|b_{k,j}+b_{i,k}|b_{k,j}|}{2}&\text{otherwise}\,.\end{cases}$
  • The cluster $\mathbf{x}^{\prime}=(x_{1}^{\prime},\ldots,x_{n}^{\prime})$ is given by $x_{i}^{\prime}=x_{i}$ if $i\neq k$, and
    $x_{k}^{\prime}=\frac{1}{(y_{k}\oplus 1)x_{k}}\left(y_{k}\prod_{B_{k,j}>0}x_{j}^{B_{k,j}}+\prod_{B_{k,j}<0}x_{j}^{-B_{k,j}}\right)\,.$
  • The coefficient tuple $\mathbf{y}^{\prime}$ is given by
    $y_{j}^{\prime}=\begin{cases}y_{j}^{-1}&i=k\,,\\ y_{j}y_{k}^{\max(0,b_{kj})}(y_{k}\oplus 1)^{-b_{kj}}&\text{otherwise}\,.\end{cases}$
@@
The matrix $B^{\prime}=\mu_{k}(B(T_{0}))$ is in fact the adjacency matrix of the triangulation obtained by flipping the edge $\tau_{k}$, and the set $\mu_{k}(\textbf{x}_{0})$ is again a cluster. Let $\mathcal{X}$ be the set of cluster variables: this is, all variables contained in all possible clusters obtained by sequences of mutations starting from the initial seed $\textbf{x}_{0}$. The cluster algebra with principal coefficients associated to the surface with marked points $(\Sigma,\mathbb{M})$ is the algebra $\mathcal{A}:=\mathbb{Q}\mathbb{P}[\mathcal{X}]$. It is independent of the chosen triangulation $T_{0}$, so it really is associated to the surface $(\Sigma,\mathbb{M})$. The cluster algebra $\mathcal{A}$ has principal coefficients (at $t_{0}$) when $\mathbf{y}_{0}=(y_{1},\ldots,y_{n})$, and this is denoted by $\mathcal{A}_{\bullet}(T_{0})$.
There is a bijection between clusters of $\mathcal{A}$ and tagged triangulations of $(\Sigma,\mathbb{M})$, as well as between cluster variables and tagged arcs in $\Sigma$. With this in mind, we will also write $y_{\tau_{i}}$ for $y_{i}$. The tagging of arcs introduced in [4] allows to fully model the mutation of all cluster variables associated to any initial tagged triangulation, since the self-folded side of a self-folded triangle cannot be flipped in the usual way.
@@ 2 Perfect matchings
We will consider only finite and loopless graphs. Let $\mathcal{G}$ be a graph with vertex set $V=\{v_{1},\ldots,v_{n}\}$ and edge set $E$. Denote by $E(i,j)$ the set of edges between $v_{i}$ and $v_{j}$. Write $s(G)$ for the underlying simple graph of $G$, and $m(i,j)=|E(i,j)|$. Of course, if $G$ is simple then $s(G)=G$ and $m(i,j)\in\{0,1\}$ for all $1\leq i,j\leq n$. If $\overrightarrow{G}$ is an orientation of $G$, then $E(i,j)$ will also denote the set of arrows with tail $v_{i}$ and head $v_{j}$.
A perfect matching of $G$ is a subset $M\subseteq E$ such that each $v\in V$ is an endpoint of exactly one member of $M$. Thus, $|M|=|V|/2$, so a graph that admits a perfect has an even number of vertices. Call such a graph even. Denote by $\operatorname{Match}(G)$ the set of perfect matchings of $G$, and let $\Phi(G):=|\operatorname{Match}(G)|.$ We will use a simple but useful device to keep track of all perfect matchings of $G$, the so-called method of variables, as in [7].
@@
(Weighted adjacency matrix) Let $G$ be a graph with vertex set $V=\{v_{1},\ldots,v_{n}\}$. For each edge $e$ of $s(G)$, consider the formal variable $z_{e}$, and let $a_{i,j}:=m(i,j)z_{e}$. The matrix $A_{G}(z):=(a_{i,j})$ is called the weighted adjacency matrix of $G$. Furthermore, if $\overrightarrow{G}$ is an orientation of $G$ with no oriented 2-cycles, write
$a^{\prime}_{i,j}:=\epsilon(i,j)m(i,j)z_{e}\,,\quad\text{where}\quad\epsilon(i,j):=\begin{cases}1&E(i,j)\neq\emptyset\,,\\ -1&E(j,i)\neq\emptyset\,,\\ 0&\text{ otherwise}\,.\end{cases}$
$A_{\overrightarrow{G}}(z):=(a^{\prime}_{i,j})$ is the weighted skew-symmetric adjacency matrix of $\overrightarrow{G}$.
@@
The matrices of definition 2.1 are defined up to a relabeling of the vertex set of $G$, and thus a simultaneous permutation of rows and columns. By setting $z_{e}=1$, we recover the usual adjacency matrices of $G$ and $\overrightarrow{G}$, so we denote them by $A_{G}(1)$ and $A_{\overrightarrow{G}}(1)$, respectively.
@@
$\phi_{G}:=\sum_{\{e_{1},\ldots,e_{k}\}\in\operatorname{Match}(G)}z_{e_{1}}\cdots z_{e_{k}}\,.$
@@
The term $c_{P}$ is independent of the order of labeling of each part in $P$. Indeed, switching the roles of $i_{k}$ and $j_{k}$ changes the sign of $\sigma_{P}$ by 1-1, but $c_{j_{k},i_{k}}=-c_{i_{k},j_{k}}$. Thus, $c_{P}$ depends only on the partition $P$. Note that such a partition corresponds to a perfect matching of $K_{2n}$, the complete graph in $2n$ vertices, and that $c_{P}\neq 0$ if and only if such a potential matching corresponds to an actual matching of $G$.
@@
Let $G$ be an even graph, and let $\overrightarrow{G}$ be an arbitrary orientation of $G$. The sum of the absolute value of the coefficients of $\operatorname{pff}(A_{\overrightarrow{G}}(z))$ is $\Phi(G)$.
@@
Take the vertex set $V$ of $G$ to be $\{1,\ldots,n\}$, and consider the simple graph $s(G)$ with edge set $E^{\prime}$. Let $k=n/2$. Each partition $P=\{\{i_{l},j_{l}\}\mid 1\leq l\leq k\}$ as in definition 2.3 corresponds to a possible set of independent edges in the simple graph $s(G)$. For $C=A_{\overrightarrow{G}}(z)$, the associated term $c_{P}$ is
@@
As observed by Tutte in [11], the previous argument and Lemma 2.1 imply that a graph $G$ admits a perfect matching if and only if there is some orientation $\overrightarrow{G}$ for which $\det(A_{\overrightarrow{G}}(z))\neq 0$. Call the polynomial $\operatorname{pff}(A_{\overrightarrow{G}}(z))$ the pfaffian of $\overrightarrow{G}$, and denote it by $\operatorname{pff}(\overrightarrow{G})$. Note that this polynomial is defined only up to a sign that depends on the chosen labeling of the vertices of $G$.
@@
Let $\overrightarrow{G}$ be an oriented graph. Then $|\operatorname{pff}(A_{\overrightarrow{G}}(1))|\leq\Phi(G)$.
An orientation $\overrightarrow{G}$ of $G$ for which the equality is attained in the previous corollary is a pfaffian orientation. Likewise, a graph is pfaffian if it admits a pfaffian orientation. For such an orientation, the pfaffian of $\overrightarrow{G}$ is a polynomial whose coefficients all have the same sign. Indeed, in such a case $\operatorname{pff}(\overrightarrow{G})=\pm\phi_{G}$. We cite a useful characterization of pfaffian orientations. Call a cycle $C$ of a graph $G$ nice if GV(C)G-V(C) admits a perfect matching. An oriented even cycle is oddly oriented if it has an odd number of arrows that agree with any of its two opposite cyclic orientations.
@@
([7], Theorem 8.3.2) Let $G$ be an even graph and $\overrightarrow{G}$ an orientation of $G$. Then $\overrightarrow{G}$ is pfaffian if and only if each even nice cycle of $G$ is oddly oriented.
Furthermore, the following is classic. A proof can be found in [7], Theorem 8.3.4.
@@
Let $\overrightarrow{G}$ be bipartite with bipartition $V=U\cup W$ such that $|U|=|W|$ (if $\overrightarrow{G}$ has a perfect matching, this is the case). The weighted skew symmetric adjacency matrix of $\overrightarrow{G}$ takes the form
$A_{\overrightarrow{G}}(z)=\begin{pmatrix}0&B_{\overrightarrow{G}}(z)\\ -B_{\overrightarrow{G}}(z)^{T}&0\end{pmatrix}\,$, (2.1)
for any labeling $V=\{u_{1},\ldots,u_{k},w_{1},\ldots,w_{l}\}$ such that $u_{i}\in U$ and $w_{j}\in W$ for $1\leq i\leq k$, $1\leq j\leq l$. $B_{\overrightarrow{G}}(x)$ is called the weighted biadjacency matrix of $\overrightarrow{G}$. The following lemma is given as an exercise in [6] (though there seems to be a sign missing therein). We include a proof.
@@
Let $B$ be a square matrix of size $n$. If
$A=\begin{pmatrix}0&B\\ -B^{T}&0\end{pmatrix}\,$,
then $\operatorname{pff}(A)=(-1)^{n(n-1)/2}\det B$.
@@
Let $V=\{1,\ldots,2n\}$, $U=\{1,\ldots,n\}$ and $W=\{n+1,\ldots,2n\}$. Fix some partition $P=\{\{i_{1},j_{1}\},\ldots,\{i_{n},j_{n}\}\}$. By remark 2.2, we can assume that $j_{i}=n+i$ for each $1\leq i\leq n$, so
@@
Note that $i_{1},\ldots,i_{n}\leq n<n+1=j_{1}$, so the indices $i_{l}$ amount to n1n-1 inversions with respect to $j_{1}$. Similarly, n2n-2 of the $i_{l}$ give n2n-2 inversions with $j_{2}$, and so on. This gives a total of n(n1)/2n(n-1)/2 inversions. Thus,
@@
Thus, $b_{P}=(-1)^{n(n-1)/2}\operatorname{sgn}(\sigma_{P}^{\prime})b_{i_{1},n+1}\cdots b_{i_{n},2n}$. Summing over all $P$ we obtain the result. ∎